# PaTAS: A Parallel System for Trust Propagation in Neural Networks Using Subjective Logic

### v2

@@ Abstract
Trustworthiness has become a key requirement for the deployment of artificial intelligence systems in safety-critical applications. Conventional evaluation metrics such as accuracy and precision fail to capture uncertainty or the reliability of model predictions, particularly under adversarial or degraded conditions. This paper introduces the Parallel Trust Assessment System (PaTAS), a framework for modeling and propagating trust in neural networks using Subjective Logic (SL). PaTAS operates in parallel with standard neural computation through Trust Nodes and Trust Functions that propagate input, parameter, and activation trust across the network. The framework defines a Parameter Trust Update mechanism to refine parameter reliability during training and an Inference-Path Trust Assessment (IPTA) method to compute instance-specific trust at inference. Experiments on real-world and adversarial datasets demonstrate that PaTAS produces interpretable, symmetric, and convergent trust estimates that complement accuracy and expose reliability gaps in poisoned, biased, or uncertain data scenarios. The results show that PaTAS effectively distinguishes between benign and adversarial inputs and identifies cases where model confidence diverges from actual reliability. By enabling transparent and quantifiable trust reasoning within neural architectures, PaTAS provides a principled foundation for evaluating model reliability across the AI lifecycle.
@@ I Introduction
Artificial Intelligence (AI) systems, particularly deep neural networks, are increasingly used in critical domains such as healthcare, autonomous driving, and cybersecurity. While highly performant, they often operate as black boxes that offer limited insight into the reliability of their outputs. This opacity becomes critical under adversarial, uncertain, or degraded input conditions. Conventional metrics such as accuracy or precision do not capture uncertainty or trust, leaving decision-makers without indicators of when to rely on model outputs. Even confidence estimates rarely account for dataset bias, label noise, or adversarial corruption. For instance, a diagnostic model may report 95% confidence in a classification even when trained on mislabeled data. These challenges underscore the need for reliability measures that go beyond predictive confidence and that consider the quality and provenance of inputs, the reliability of training data, and the stability of learned parameters.
@@
As AI systems become integral to real-world applications, trustworthiness has emerged as a fundamental requirement for ensuring reliability, safety, and alignment with human values. According to the High-Level Expert Group on AI [1], a trustworthy system must be lawful, ethical, and robust. In technical terms, this translates into measurable properties such as accuracy, robustness, fairness, and explainability [2]. Embedding these properties across the AI pipeline, from data collection to deployment, is essential for preventing failures caused by poor data quality, bias, or unstable training. In this work, we adopt a property-based view of trustworthiness that focuses on quantifiable technical dimensions which can be modeled and propagated within neural networks. Specifically, we study how trust in the dataset, the query input, and the learned parameters (assessed for the same property) jointly determine the reliability of model predictions.
@@
These issues become even more pronounced in transfer learning scenarios, where knowledge is reused across domains or tasks. Transfer learning introduces additional challenges related to both the transferability and reliability of knowledge. Recent studies highlight the need for frameworks that assess whether transferred knowledge remains robust, fair, and secure across varying data distributions [3]. Such findings reinforce the broader need for a unified approach to measuring and propagating trust in complex neural architectures.
@@
Despite this growing awareness, most models still provide limited mechanisms for representing confidence. Neural networks are typically trained on clean, unambiguous data and thus lack exposure to uncertain or borderline cases. As a result, predicted probabilities tend to reflect similarity to seen patterns rather than genuine uncertainty, and they often assume clean, in-distribution inputs. These assumptions are rarely satisfied in real-world or adversarial settings, where data quality, distributional shifts, and parameter stability all influence prediction reliability. Consequently, existing approaches lack mechanisms to evaluate how uncertainty and trust in inputs, activations, and parameters jointly affect model behavior. This motivates the development of tools that can explicitly reason about trust across the entire model lifecycle.
Problem
@@
Existing methods for trust and uncertainty estimation in neural networks face several key limitations:
@@
    Limited holistic propagation: Uncertainty quantification methods typically assess reliability only at the output layer. Even attribution-based methods (e.g., SmoothGrad [4]) fail to capture how input, intermediate, and parameter trust jointly influence the output reliability.
@@
    Lack of interpretability: Few models yield trust estimates that are both faithful to the model’s reasoning and understandable to end users.
These limitations hinder reliable trust assessments, particularly in safety-critical or adversarial contexts.
@@
To address these challenges, we propose the Parallel Trust Assessment System (PaTAS), a framework for modeling and propagating trust in neural networks using Subjective Logic (SL). The main contributions are summarized as follows:
@@
    Parameter Trust Update: We design an algorithm that refines trust in learned parameters using gradient behavior, input trust, and label trust, aligning parameter reliability with the learning dynamics.
@@
The remainder of this paper is organized as follows: Section II introduces background on Subjective Logic and dataset trust assessment, Section III reviews related work, Section IV formalizes trust propagation in neural networks, Section V details the PaTAS architecture, Section VI presents experiments, Section VII discusses implications, and Section VIII concludes the paper.
@@ II-A Subjective Logic Fundamentals
Reasoning about trust in AI systems requires handling partial, conflicting, or missing evidence. Classical probability theory models aleatoric uncertainty but cannot represent ignorance or contradictory information, which often arises from noisy or biased data. Subjective Logic (SL) [5] extends Dempster–Shafer theory [6, 7] to capture these conditions by representing opinions rather than probabilities. An SL opinion expresses belief, disbelief, and uncertainty as separate components, thereby distinguishing between uncertainty due to lack of knowledge (epistemic uncertainty) and uncertainty inherent to the environment itself (aleatoric uncertainty).
A subjective opinion denoted by $\omega_{X}^{A}$ expresses the belief of an agent $A$ about states of a variable $X$ which takes its values from a domain $\mathbb{X}$ (i.e., a state space). A special case of a subjective opinion is a subjective binomial opinion where $card(\mathbb{X})=2$. For a binary variable $X\in\mathbb{X}=\{x,\overline{x}\}$, a binomial opinion is expressed as a quadruple:
@@
Binomial opinions can be derived from evidence through various quantification models. Let $r_{x}$ and $s_{x}$ represent the amount of positive and negative evidence. Three common quantification approaches are:
@@
A subjective opinion is meaningful only within a specific context or property under evaluation (e.g., accuracy, bias, or other trust aspects). In this work, trust is represented as a subjective binomial opinion $(t,d,u)$, where $t$ denotes trust (belief), $d$ distrust (disbelief), and $u$ uncertainty. These fundamentals define how trust is represented and interpreted as subjective opinions. To make them operational, Subjective Logic provides operators for combining, revising, and discounting opinions, which we introduce next.
@@ II-B Subjective Logic Operators
Subjective Logic (SL) provides key reasoning operators for combining and propagating opinions, including trust discounting, fusion, and inferential operators [9, 10, 5].
@@
The specific fusion operator depends on the relationship between the sources. SL defines several variants—such as consensus, averaging, weighting, and cumulative fusion—each representing a distinct way of aggregating evidence.
Example. Suppose two temperature sensors estimate whether the room temperature exceeds $25^{\circ}$C. If both sensors are of the same type and environment, their evidence is correlated and averaging or weighted fusion is appropriate. If they are of different designs (e.g., infrared and contact-based) and thus independent, cumulative fusion is more suitable.
@@
Let $A$ have a referral trust $\omega^{A}_{B}$ in another agent $B$, who holds an opinion $\omega^{B}_{X}$ on variable $X$. The trust-discounted opinion of $A$ derived from $B$ is:
@@
Example. Continuing the sensor case, assume one sensor occasionally drifts. Agent $A$ assigns a referral trust to that sensor $B$, discounting its opinion to reduce belief mass and increase uncertainty. This ensures fusion incorporates both evidence and source reliability.
@@
Inferential operators generalize Bayesian reasoning by enabling opinion propagation through conditional relationships. If $A$ reasons about a relation $X\Rightarrow Y$ with conditional opinion $\omega^{A}_{Y\mid X}$, belief ($b$), disbelief ($d$), and uncertainty ($u$) describe how $X$ supports $Y$. The main operators are:
@@
Example. For the relation “if it rains ($X$), Bob carries an umbrella (YY),” agent $A$ holds a conditional opinion $\omega^{A}_{Y\mid X}=(b,d,u)$. Given an opinion on rain, deductive inference yields an opinion about umbrella use; conversely, abduction infers the likelihood of rain from umbrella observations.
A summary of all operators and symbols used in this work appears in Appendix A.
@@
A Subjective Trust Network (STN) [12, 13] models trust relationships as subjective opinions propagated through referral chains using trust discounting and fusion. Such extensions of SL have been applied to trust propagation, opinion dynamics, and source reliability analysis [14, 15, 16, 17]. In practical domains such as automotive systems, SL has been used to combine evidence from multiple misbehavior detection mechanisms in vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication, improving resilience against insider attacks [18, 19].
@@ II-C Trustworthiness in the dataset
The quality and structure of the training dataset are essential for determining the performance, robustness, and fairness of machine learning models. Common issues such as sampling biases, mislabeled instances, or a lack of diversity in the data can degrade learned representations and hinder generalization, thereby reducing the reliability of the model’s outputs [20]. Recent studies further show that even small corrections in large-scale datasets can significantly alter model performance. For example, on CIFAR-10 with corrected labels, VGG11 outperforms VGG19 once the prevalence of mislabeled test examples increases by just 5%, illustrating how pervasive label errors can destabilize benchmark conclusions and affect model selection [20]. Thus, evaluating the trustworthiness of training data is a critical step in assessing the trustworthiness of an AI system.
@@
  • Dataset level: Captures global properties such as class imbalance or sampling bias. These factors affect the overall distribution and may harm generalization.
  • Instance level: Addresses local anomalies like mislabeled or corrupted data points. Individual instances may be unreliable due to noise, annotation errors, or improper data collection.
  • Input feature level: Trust assessments at this level capture fine-grained variations within the input vector. For instance, in a data poisoning scenario, an adversarial patch affecting only one specific pixel may render part of an image untrustworthy [21]. Similarly, when input feature values are sourced from heterogeneous systems, some features may be inherently more reliable than others, leading to variable trust across features.
@@ III Related Work
Quantifying trust and uncertainty in neural networks has become a central research topic, particularly in safety-critical domains. Incorrect yet confident predictions can have severe consequences. For instance, in data poisoning, if a model is trained and evaluated on corrupted or mislabeled samples, it may still report high accuracy, precision, or recall, creating an illusion of reliability. Standard metrics therefore fail to capture subtleties of model trustworthiness. As a result, diverse frameworks have been proposed to model and propagate uncertainty and trust, yet significant limitations remain.
@@ III-A Uncertainty Quantification in Neural Networks
Uncertainty quantification (UQ) aims to estimate predictive reliability by modeling uncertainty at different levels, typically epistemic (model-based) and aleatoric (data-based). Foundational methods include Bayesian neural networks [22, 23, 24], Monte Carlo dropout [25], and ensembles [26]. Dropout provides an efficient Bayesian approximation [25], while Bayes by Backprop [23] learns weight distributions that enhance generalization and exploration in reinforcement learning. Extensions with latent variables explicitly decompose predictive uncertainty into epistemic and aleatoric components [24], improving decision-making in active and reinforcement learning through risk-sensitive criteria.
Despite these advances, estimating and calibrating both forms of uncertainty in complex models remains challenging [27]. Studies report frequent over- or under-confidence, with uncertainty estimates often degrading under real-world conditions [28, 29, 30]. In medical AI, uncalibrated confidence has been linked to critical misjudgments [28]. Even well-performing models can produce unreliable confidence scores under dataset shifts [30]. Consequently, researchers have explored post-hoc calibration, such as temperature scaling [31], which adjusts output probabilities to better align predicted and observed frequencies. However, these approaches operate only at the output layer and assume clean data, which can still yield misleading confidence when inputs or training data are corrupted.

### v3

# PaTAS: A Framework for Trust Propagation in Neural Networks Using Subjective Logic

@@ Abstract
Trustworthiness has become a key requirement for the deployment of artificial intelligence systems in safety-critical applications. Conventional evaluation metrics, such as accuracy and precision, fail to appropriately capture uncertainty or the reliability of model predictions, particularly under adversarial or degraded conditions. This paper introduces the Parallel Trust Assessment System (PaTAS), a framework for modeling and propagating trust in neural networks using Subjective Logic (SL). PaTAS operates in parallel with standard neural computation through Trust Nodes and Trust Functions that propagate input, parameter, and activation trust across the network. The framework defines a Parameter Trust Update mechanism to refine parameter reliability during training and an Inference-Path Trust Assessment (IPTA) method to compute instance-specific trust at inference. Experiments on real-world and adversarial datasets demonstrate that PaTAS produces interpretable, symmetric, and convergent trust estimates that complement accuracy and expose reliability gaps in poisoned, biased, or uncertain data scenarios. The results show that PaTAS effectively distinguishes between benign and adversarial inputs and identifies cases where model confidence diverges from actual reliability. By enabling transparent and quantifiable trust reasoning within neural architectures, PaTAS provides a foundation for evaluating model reliability across the AI lifecycle.
@@ I Introduction
Artificial Intelligence (AI) systems, particularly Neural Networks (NNs) [1] are being employed in critical sectors such as healthcare to interpret clinical images [2] or in autonomous driving to recognize physical elements in traffic images [3]. While highly performant, they often operate as black boxes that offer limited insight into the reliability of their outputs. This opacity becomes critical under adversarial, uncertain, or degraded input conditions. Conventional performance metrics such as accuracy or precision do not capture uncertainty and reliability, leaving decision-makers with misleading indicators of a model’s outputs quality [4]. Even confidence estimates rarely account for dataset bias, label noise, or adversarial corruption. For instance, a diagnostic model may report 95% confidence in a classification task while being evaluated on mislabeled data. These challenges underscore the need for reliability measures that go beyond predictive confidence and that consider the quality and provenance of inputs, the reliability of training data, and the stability of learned parameters.
@@
As AI systems become integral to critical real-world applications, trustworthiness has emerged as a fundamental requirement for ensuring reliability, safety, and alignment with human values. According to the High-Level Expert Group on AI [5], a trustworthy system must be lawful, ethical, and robust. In technical terms, this translates into measurable properties such as accuracy, robustness, fairness, and explainability [6]. Embedding these properties across the AI pipeline, from data collection to deployment, is essential for preventing failures caused by poor data quality, bias, or unstable training. In this work, we adopt a property-based view of trustworthiness, where a property refers to a specific technical aspect such as accuracy or bias. We focus on quantifiable technical dimensions that can be modeled and propagated within neural networks. Specifically, we study how trustworthiness in the dataset, the query input, and the learned parameters (assessed for the same property) jointly determine the reliability of model predictions.
@@
These issues become even more pronounced in transfer learning scenarios, where knowledge is reused across domains or tasks. Transfer learning introduces additional challenges related to both the transferability and reliability of knowledge. A recent study [7] highlights the need for frameworks that assess whether transferred knowledge remains robust, fair, and secure across varying data distributions. These findings emphasize the need for a uniform approach for evaluating and propagating trustworthiness in complex neural networks.
@@
Despite this growing awareness, most models still provide limited mechanisms for representing confidence in their outputs. Neural networks are typically trained with the assumption of clean, unambiguous data labels and thus lack in the capability to generalize across previously unseen borderline cases. As a result, predicted probabilities tend to reflect similarity to seen patterns rather than genuine uncertainty, and they often assume clean, in-distribution inference inputs. These assumptions are rarely satisfied in real-world or adversarial settings, where data quality, distributional shifts, and parameter stability all influence prediction reliability. Consequently, existing approaches lack mechanisms to evaluate how uncertainty and trustworthiness in training inputs, activations, and parameters jointly affect model behavior. Even attribution-based methods (like SmoothGrad [8]) fail to capture how training input, intermediate, and parameter trustworthiness jointly influence the output reliability. This gap motivates the development of tools that can explicitly reason about trustworthiness across the entire model lifecycle.
Problem Statement
@@
In summary, existing methods for trustworthiness and uncertainty estimation in neural networks face several key limitations:
@@
    Limited holistic propagation: Uncertainty quantification methods typically assess reliability only at the output layer.
@@
    Lack of interpretability: Few models yield trustworthiness estimates that are both faithful to the model’s reasoning and understandable to end users.
These limitations hinder reliable trustworthiness assessments, particularly in safety-critical or adversarial contexts.
@@
To address these challenges, we propose the Parallel trustworthiness Assessment System (PaTAS), a framework for modeling and propagating trustworthiness in NNs using Subjective Logic (SL). The main contributions are summarized as follows:
@@
    Parameter Trust Update: We design an algorithm that determines trust in learned parameters using gradient values, input trust, and label trust, and aligning parameter reliability with the learning dynamics.
@@
The remainder of this paper is organized as follows: Section II introduces background on Subjective Logic and dataset trustworthiness assessment, Section III reviews related work, Section IV formalizes trust propagation in neural networks, Section V details the PaTAS architecture, Section VI presents experiments, Section VII discusses implications, and Section VIII concludes the paper and provides an outlook on future work.
@@ II Background
Trust assessment in NNs requires a formalization that may represent ambiguity, inadequate evidence, and source reliability. Section II-A summarizes the principles of Subjective Logic (SL) and the trust opinion representation, which encodes trust, distrust, and uncertainty. Section II-B covers the primary SL reasoning operators and their application in Subjective Trust Networks. Finally, we explain in Section II-C how these concepts manifest as measures of dataset trustworthiness.
@@ II-A Subjective Logic Fundamentals
Reasoning about trust in AI systems requires handling partial, conflicting, or missing evidence. Classical probability theory models aleatoric uncertainty but cannot represent missing knowledge (ignorance) or contradictory information, which often arises from noisy or biased data. Subjective Logic [9] extends Dempster–Shafer theory [10, 11] to capture these conditions by representing trust as SL opinions rather than probabilities. A subjective opinion expresses beliefs and uncertainty as separate components, thereby distinguishing between uncertainty due to lack of knowledge (epistemic uncertainty) and uncertainty inherent to the environment itself (aleatoric uncertainty).
A subjective opinion denoted by $\omega_{X}^{A}$ expresses the beliefs of an agent $A$ (e.g., a sensor, a human, or an external observer of a process) about states of a variable $X$ which takes its values from a domain $\mathbb{X}$ (i.e., a state space). A special case of a subjective opinion is a subjective binomial opinion where $card(\mathbb{X})=2$. For a binary variable $X\in\mathbb{X}=\{x,\overline{x}\}$, a binomial opinion is expressed as a quadruple:
@@
This projection reduces the richer subjective opinion to an equivalent classical probability, enabling compatibility with standard probabilistic reasoning.
Binomial opinions can be derived from evidence through various quantification models. Let $r_{x}$ and $s_{x}$ represent the amount of positive and negative evidence. Positive evidence $r_{x}$ captures observations that increase confidence in the truth of $x$, while negative evidence $s_{x}$ captures observations that support its falsehood. Three common quantification approaches are:
@@
    where a fixed uncertainty distributes the remaining mass.
A subjective opinion is meaningful only within a specific context or property under evaluation (e.g., accuracy, bias, or other trust-related aspects). In this work, trust is represented as a subjective binomial opinion $(t,d,u)$, where $t$ denotes trust (belief), $d$ distrust (disbelief), and $u$ uncertainty. These fundamentals define how trust is represented and interpreted as subjective opinions. To make them operational, Subjective Logic provides operators for combining, revising, and discounting opinions, which we introduce next.
@@ II-B Subjective Logic Operators
SL provides key reasoning operators for combining and propagating opinions, including trust discounting, fusion, and inferential operators [13, 14, 9].
@@
The specific fusion operator depends on the relationship between the sources. SL defines several variants such as consensus, averaging, weighting, and cumulative fusion, each representing a distinct way of aggregating evidence.
Example. Suppose two temperature sensors estimate whether the room temperature exceeds $25^{\circ}$C. If both sensors are of the same type and installed in the same place, their evidence is correlated and averaging or weighted fusion is appropriate. If they have different measurement strategies (e.g., infrared and contact-based), cumulative fusion is more suitable. The information from each independent sensor is treated as an additional, non-redundant contribution so that the combined evidence grows with each agreeing source.
@@
Let $A$ have a referral trust $\omega^{A}_{B}$ in another agent $B$, who holds an opinion $\omega^{B}_{X}$ on variable $X$. The trust-discounted opinion of $A$ derived from $B$’s opinion is:
@@
Example. Consider a monitoring system where agent $A$ receives readings from a temperature sensor $B$. The sensor usually works well but is known to drift at times, so $A$ does not fully trust it. When $B$ reports that the temperature is above a safety threshold, $A$ discounts this opinion by reducing its strength and increasing uncertainty. This illustrates the principle of trust discounting: evidence from a partially reliable source is treated cautiously.
@@
Inferential operators generalize Bayesian reasoning by enabling opinion propagation through conditional relationships. Let $A$ be an agent reasoning about a variable $X$ and its potential implications for another variable $Y$. Suppose $A$ holds opinions $\omega^{A}_{Y\mid X}$ on the relationship $X\Rightarrow Y$. The main operators are:
@@
Example. For the relation “if it rains ($X$), then Bob carries an umbrella (YY),” agent $A$ holds a conditional opinion $\omega^{A}_{Y\mid X}=(b,d,u)$. Given an opinion on rain, deductive inference yields an opinion about umbrella use; conversely, abduction infers the likelihood of rain from umbrella observations.
A summary of all operators used in this work and their symbols appears in Appendix A.
@@
A Subjective Trust Network (STN) [16, 17] models trust relationships as subjective opinions propagated through referral chains using trust discounting and fusion. Such extensions of SL have been applied to trust propagation, opinion dynamics, and source reliability analysis [18, 19, 20, 21].
@@ II-C Trustworthiness in the dataset
Focusing back on machine learning, the quality and structure of the training dataset are essential for determining the performance, robustness, and fairness of machine learning models. Common issues such as sampling biases, mislabeled instances, or lack of diversity in the data can degrade learned representations and hinder generalization, thereby reducing the reliability of the model’s outputs [22]. A recent study on dataset quality shows that even small proportions of mislabeled samples can substantially shift model rankings; for instance, on CIFAR-10, VGG11 trained on clean data can outperform VGG19 once the fraction of erroneous labels reaches about 5%, illustrating how sensitive benchmark conclusions are to label errors [23]. Thus, evaluating the trustworthiness of training data is a critical step in assessing the trustworthiness of an AI system.
@@
  • Dataset level: Captures global explanations of misbehavior such as class imbalance or sampling bias. These factors affect the overall distribution and may harm generalization.
  • Instance level: Captures local explanations of anomalies like mislabeled or corrupted data points. Individual instances may be unreliable due to noise, annotation errors, or improper data collection.
  • Input feature level: Captures fine-grained variations within the input vector. For instance, in a data poisoning scenario, an adversarial patch affecting only one specific pixel may render part of an image untrustworthy [24]. Similarly, when input feature values are sourced from heterogeneous systems, some features may be inherently more reliable than others, leading to variable trust across features.
@@ III Related Work
Quantifying trust and uncertainty in neural networks has become a central research topic, particularly in safety-critical domains. Incorrect yet confident predictions can have severe consequences. For example, data poisoning occurs when both the training and test datasets contain systematically mislabeled samples. However, the model may obtain good accuracy, precision, and recall on the poisoned test set. This observation creates the illusion of dependability despite having learned harmful or erroneous patterns. Standard metrics therefore fail to capture subtleties of model trustworthiness. As a result, diverse frameworks have been proposed to model and propagate uncertainty and trust, yet significant limitations remain as we will point out next.
@@ III-A Uncertainty Quantification in Neural Networks
Uncertainty quantification (UQ) aims to estimate predictive reliability by modeling uncertainty at different levels, typically epistemic (model-based) and aleatoric (data-based) [25]. Foundational methods include Bayesian neural networks [26, 27, 28], Monte Carlo dropout [29], and ensembles [30]. Dropout provides an efficient Bayesian approximation [29], while Bayes by Backprop [27] learns weight distributions that enhance generalization and exploration in reinforcement learning. Extensions with latent variables explicitly decompose predictive uncertainty into epistemic and aleatoric components [28], improving decision-making in active and reinforcement learning through risk-sensitive criteria.
Despite these advances, estimating and calibrating both forms of uncertainty in complex models remains challenging [31]. Studies report frequent over- or under-confidence, with uncertainty estimates often degrading under real-world conditions [32, 33, 34]. In medical AI, uncalibrated confidence has been linked to critical misjudgments [2, 32]. Even well-performing models can produce unreliable confidence scores under dataset shifts [34]. Consequently, researchers have explored post-hoc calibration, such as temperature scaling [35], which adjusts output probabilities to better align predicted and observed frequencies. However, these approaches operate only at the output layer and assume clean data, which can still yield misleading confidence when inputs or training data are corrupted.